OpenStudy (anonymous):

Use Euler's formula to write the given expression e^2-(pi/2)i in the form a+ib.

OpenStudy (loser66):

is it \[\huge e^{2-\frac{\pi}{2}i}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Wait a minute. Please don't leave.

hartnn (hartnn):

strange question indeed, but if its that, then i'll do \(\dfrac{e^2}{e^{i \pi/2}}\) and find e^(i pi/2) first

hartnn (hartnn):

do u know how to get e^(i pi/2) ??

hartnn (hartnn):

use same formula as i gave in last question, \(e^{i \theta}=\cos \theta+i\sin \theta\)

OpenStudy (anonymous):

Wait a minute, please don't leave.

OpenStudy (anonymous):

So I got e^2/i, now what?

OpenStudy (anonymous):

Because e^i*pi/2=i.

hartnn (hartnn):

that could be the answer but the denominator is not real, to rationalize the denominator, you can multiply and divide by 'i' and use the fact that i^2=-1 in the denominator

OpenStudy (anonymous):

I forgot to say thank you.

hartnn (hartnn):

no problem :) i'll forget to say welcome :P